OpenStudy (anonymous):

Based on the table of values below, find the slope between points where x = 1 and where x = 4. x y 1 8 3 6 4 −1 answer choices: −3 -2/3 3/2 3

OpenStudy (johnweldon1993):

I believe they want the average rate of change here between 1 and 4 \[\large \frac{f(b) - f(a)}{b - a}\] b = 4 a = 1 So what does the function equal when x= 4? From the table we see -1 What about when x = 1? From the table we have 8 So \[\large \frac{-1 - 8}{4 - 1} = \frac{-9}{3} = -3\]

OpenStudy (anonymous):

thank you so much! (:

OpenStudy (johnweldon1993):

Anytime :)